# If in a Crowdsourced Data Annotation Pipeline, a GPT-4

### v1

@@ Abstract.
Recent studies indicated GPT-4 outperforms online crowd workers in data labeling accuracy, notably workers from Amazon Mechanical Turk (MTurk). However, these studies were criticized for deviating from standard crowdsourcing practices and emphasizing individual workers’ performances over the whole data-annotation process. This paper compared GPT-4 and an ethical and well-executed MTurk pipeline, with 415 workers labeling 3,177 sentence segments from 200 scholarly articles using the CODA-19 scheme. Two worker interfaces yielded 127,080 labels, which were then used to infer the final labels through eight label-aggregation algorithms. Our evaluation showed that despite best practices, MTurk pipeline’s highest accuracy was 81.5%, whereas GPT-4 achieved 83.6%. Interestingly, when combining GPT-4’s labels with crowd labels collected via an advanced worker interface for aggregation, 2 out of the 8 algorithms achieved an even higher accuracy (87.5%, 87.0%). Further analysis suggested that, when the crowd’s and GPT-4’s labeling strengths are complementary, aggregating them could increase labeling accuracy.
@@ 1. Introduction
This paper presents a holistic investigation that compares the data labeling abilities of GPT-4 and a well-executed, ethical MTurk data-annotation process. We had 415 MTurk workers label 3,177 sentence segments in 200 scholarly papers published in 2022 or later, where each segment collected 20 labels. We used the CODA-19 dataset’s 5-class label scheme (Huang et al., 2020), categorizing sentence segments by their research aspects: Background, Purpose, Method, Finding/Contribution, and Other. We further experimented with two distinctly designed worker interfaces, recognizing the potential biases of any particular annotation interface (Rahmanian and Davis, 2014; Toomim et al., 2011). A total of 127,080 labels was gathered in the study (3,177 sentence segments $\times$× 2 interface variations $\times$× 20 labels.) [^2] We will keep the dataset offline for a minimum of one year to avoid data contamination concerns of large language models pre-trained using data on the public Internet. Access to the dataset will be email-exclusive, and we will mandate recipients not to upload it to the Internet. Before sharing the data, we will hash the worker IDs to ensure they are unique within the dataset but unrecognizable as real worker IDs. We then applied 8 label aggregation algorithms, including Majority Voting and DawidSkene, to determine the final labels and compared the labeling accuracies with GPT-4. We found that, even with the best crowdsourcing practices, MTurk’s top-performing pipeline’ accuracy of 81.5% did not surpass GPT-4’s 83.6% (p¡0.05). Interestingly, when combining GPT-4’s labels with crowd labels collected via an advanced worker interface for aggregation, 2 out of the 8 algorithms achieved a higher accuracy (87.5% with p¡0.01 and 87.0% with p¡0.01) compared to GPT-4’s standalone performance (83.6%). Further analysis suggested that, when the crowd’s and GPT-4’s labeling strengths are complementary– with crowds better at labeling the Finding/Contribution class and GPT-4 excelling in all other classes– aggregating them could further increase labeling accuracy.
@@ 4. Experimental Results
As a result, while more crowd labels improved accuracy, even aggregating 20 could not surpass GPT-4. We also noticed that in the simulations, accuracy improved for most aggregation methods as the number of MTurk workers increased, but MACE’s performance was notably inconsistent.
@@ 4.2. GPT-4 in MTurk Pipelines
To further aid in interpreting these results, we calculated 95% confidence intervals for sentence-level accuracy (N=3,177). Tables 3, 4, 6, and 7 present these accuracies alongside their respective 95% confidence intervals, highlighting the precision and reliability of our findings.
@@
MV .551 ¡.001 [.534, .569] .567 ¡.001 [.550, .584] .609 ¡.001 [.592, .626]
DawidSkene .675 ¡.001 [.659, .691] .712 ¡.001 [.696, .727] .737 ¡.001 [.722, .753]
OneCoin .797 .002 [.783, .811] .800 .006 [.786, .814] .804 .002 [.790, .818]
GLAD .734 ¡.001 [.719, .749] .731 ¡.001 [.715, .746] .756 ¡.001 [.741, .771]
M-MSR .531 ¡.001 [.513, .548] .590 ¡.001 [.572, .607] .630 ¡.001 [.614, .647]
MACE .811 .001 [.797, .824] .809 ¡.001 [.796, .823] .809 .001 [.795, .823]
Wawa .633 ¡.001 [.617, .650] .636 ¡.001 [.619, .653] .672 ¡.001 [.656, .689]
ZBS .677 ¡.001 [.661, .694] .675 ¡.001 [.659, .691] .712 ¡.001 [.696, .727]
@@ 4.3. Analyzing Cases Where Crowd Labels Enhance GPT’s Label Accuracy
Figure 7. One-Coin Dawid-Skene and MACE Flip to Correct/Incorrect Stacked Bar Chart for Advanced Interface under Exclude-By-Worker, comparing to GPT-Only results, respectively. Figure 6(a) and 6(b) show that crowd workers benefit GPT results mainly in the “Finding” category (positive X-axis), while the decline in some of GPT’s annotations across all labels at different levels (negative X-axis)
@@
Figure 7. One-Coin Dawid-Skene and MACE Flip to Correct/Incorrect Stacked Bar Chart for Advanced Interface under Exclude-By-Worker, comparing to GPT-Only results, respectively. Figure 6(a) and 6(b) show that crowd workers benefit GPT results mainly in the “Finding” category (positive X-axis), while the decline in some of GPT’s annotations across all labels at different levels (negative X-axis)
@@ Appendix A All Workers and Exclude-By-Batch Result Table
In this section, we show all the additional aggregation results for different distinct configurations tested (2 different interfaces, and 3 different cleaning strategies). Table 8 to 12 is the aggregation accuracy results on both crowd-only and crowd+gpt results for both Basic and Advanced Interfaces under different cleaning strategies. They show detailed P, R, F1, Accuracy, and Kappa for each aggregation model under different user interfaces.
@@ Appendix C Prompt
[Finding]: Text segments answer one or more of these questions: What did the researchers find out?, Did the proposed methods work?, Did the thing behave as the researchers expected?
[Other]: Text fragments that do NOT fit into any of the four categories above. Text fragments that are NOT part of the article. Text fragments that are NOT in English. Text fragments that contains ONLY reference marks (e.g., ”[1,2,3,4,5”) or ONLY dates (e.g., ”April 20, 2008”). Captions for figures and tables (e.g. ”Figure 1: Experimental Result of …”, or ”Table 1: The Typical Symptoms of …”) Formatting errors. I really don’t know or I’m not sure.

### v2

@@ Abstract.
Recent studies indicated GPT-4 outperforms online crowd workers in data labeling accuracy, notably workers from Amazon Mechanical Turk (MTurk). However, these studies were criticized for deviating from standard crowdsourcing practices and emphasizing individual workers’ performances over the whole data-annotation process. This paper compared GPT-4 and an ethical and well-executed MTurk pipeline, with 415 workers labeling 3,177 sentence segments from 200 scholarly articles using the CODA-19 scheme. Two worker interfaces yielded 127,080 labels, which were then used to infer the final labels through eight label-aggregation algorithms. Our evaluation showed that despite best practices, MTurk pipeline’s highest accuracy was 81.5%, whereas GPT-4 achieved 83.6%. Interestingly, when combining GPT-4’s labels with crowd labels collected via an advanced worker interface for aggregation, 2 out of the 8 algorithms achieved an even higher accuracy (87.5%, 87.0%). Further analysis suggested that, when the crowd’s and GPT-4’s labeling strengths are complementary, aggregating them could increase labeling accuracy. [^1] Minor errors were corrected on June 26, 2024. These errors did not impact the conclusion of the paper. The information in this arXiv paper has been corrected, and the corrigendum is attached at the end (Appendix E).
@@ 1. Introduction
This paper presents a holistic investigation that compares the data labeling abilities of GPT-4 and a well-executed, ethical MTurk data-annotation process. We had 415 MTurk workers label 3,177 sentence segments in 200 scholarly papers published in 2022 or later, where each segment collected 20 labels. We used the CODA-19 dataset’s 5-class label scheme (Huang et al., 2020), categorizing sentence segments by their research aspects: Background, Purpose, Method, Finding/Contribution, and Other. We further experimented with two distinctly designed worker interfaces, recognizing the potential biases of any particular annotation interface (Rahmanian and Davis, 2014; Toomim et al., 2011). A total of 127,080 labels was gathered in the study (3,177 sentence segments $\times$× 2 interface variations $\times$× 20 labels.) [^3] We will keep the dataset offline for a minimum of one year to avoid data contamination concerns of large language models pre-trained using data on the public Internet. Access to the dataset can be obtained by submitting a request to the first author. Before sharing the data, we will hash the worker IDs to ensure they are unique within the dataset but unrecognizable as real worker IDs. We then applied 8 label aggregation algorithms, including Majority Voting and DawidSkene, to determine the final labels and compared the labeling accuracies with GPT-4. We found that, even with the best crowdsourcing practices, MTurk’s top-performing pipeline’ accuracy of 81.5% did not surpass GPT-4’s 83.6% (p¡0.05). Interestingly, when combining GPT-4’s labels with crowd labels collected via an advanced worker interface for aggregation, 2 out of the 8 algorithms achieved a higher accuracy (87.5% with p¡0.01 and 87.0% with p¡0.01) compared to GPT-4’s standalone performance (83.6%). Further analysis suggested that, when the crowd’s and GPT-4’s labeling strengths are complementary– with crowds better at labeling the Finding/Contribution class and GPT-4 excelling in all other classes– aggregating them could further increase labeling accuracy.
@@ 4. Experimental Results
As a result, while more crowd labels improved accuracy, even aggregating 20 could not surpass GPT-4. We also noticed that in the simulations, accuracy improved for most aggregation methods as the number of MTurk workers increased.
@@ 4.2. GPT-4 in MTurk Pipelines
To further aid in interpreting these results, we calculated 95% confidence intervals for sentence-level accuracy (N=3,177). Tables 3, 4, 14, and 7 present these accuracies alongside their respective 95% confidence intervals, highlighting the precision and reliability of our findings.
@@
MV .551 ¡.001 [.534, .569] .609 ¡.001 [.592, .626] .567 ¡.001 [.550, .584]
DawidSkene .675 ¡.001 [.659, .691] .737 ¡.001 [.722, .753] .712 ¡.001 [.696, .727]
OneCoin .797 .002 [.783, .811] .804 .002 [.790, .818] .800 .006 [.786, .814]
GLAD .734 ¡.001 [.719, .749] .756 ¡.001 [.741, .771] .731 ¡.001 [.715, .746]
M-MSR .531 ¡.001 [.513, .548] .630 ¡.001 [.614, .647] .590 ¡.001 [.572, .607]
MACE .811 .001 [.797, .824] .809 .001 [.795, .823] .809 ¡.001 [.796, .823]
Wawa .633 ¡.001 [.617, .650] .672 ¡.001 [.656, .689] .636 ¡.001 [.619, .653]
ZBS .677 ¡.001 [.661, .694] .712 ¡.001 [.696, .727] .675 ¡.001 [.659, .691]
@@ 4.3. Analyzing Cases Where Crowd Labels Enhance GPT’s Label Accuracy
Figure 7. One-Coin Dawid-Skene and MACE Flip to Correct/Incorrect Stacked Bar Chart for Advanced Interface under Exclude-By-Worker, comparing to GPT-Only results, respectively. Figure 7(a) and 7(b) show that crowd workers benefit GPT results mainly in the “Finding” category (positive X-axis), while the decline in some of GPT’s annotations across all labels at different levels (negative X-axis)
@@ Appendix A All Workers and Exclude-By-Batch Result Table
In this section, we show all the additional aggregation results for different distinct configurations tested (2 different interfaces, and 3 different cleaning strategies). Table 8 to 15 is the aggregation accuracy results on both crowd-only and crowd+gpt results for both Basic and Advanced Interfaces under different cleaning strategies. They show detailed P, R, F1, Accuracy, and Kappa for each aggregation model under different user interfaces.
@@ Appendix C Prompt
[Finding]: Text segments answer one or more of these questions: What did the researchers find out?, Did the proposed methods work?, Did the thing behave as the researchers expected? [Other]: Text fragments that do NOT fit into any of the four categories above. Text fragments that are NOT part of the article. Text fragments that are NOT in English. Text fragments that contains ONLY reference marks (e.g., ”[1,2,3,4,5”) or ONLY dates (e.g., ”April 20, 2008”). Captions for figures and tables (e.g. ”Figure 1: Experimental Result of …”, or ”Table 1: The Typical Symptoms of …”) Formatting errors. I really don’t know or I’m not sure.
@@ Appendix D All Aggregation Methods for All Workers and Exclude-By-Batch
Appendix E Corrigendum
Minor errors were corrected on June 26, 2024. These errors did not impact the conclusion of the paper. The information in this arXiv paper has been corrected. We present the corrigendum in the following:
Corrigendum to “If in a Crowdsourced Data Annotation Pipeline, A GPT-4”, by He et al., CHI ’24: Proceedings of the CHI Conference on Human Factors in Computing Systems, May 2024, Article No.: 1040, Pages 1-25
This corrigendum concerns our paper in CHI ’24: Proceedings of the CHI Conference on Human Factors in Computing Systems, May 2024, Article No.: 1040, Pages 1-25 (He et al., If in a Crowdsourced Data Annotation Pipeline, A GPT-4).
  1. (1)
    The results presented in Table 6 in the original paper should be replaced with the results in Table 14 in this document.
    In Table 6 in the original paper, the data for the eight aggregation methods listed under the ’Exclude-By-Worker’ and ’Exclude-By-Batch’ categories were mistakenly swapped during the editing process. The ‘Avg. Acc’ and ‘#workers’ rows were correct.
  2. (2)
    The results presented in Table 12 in the original paper should be replaced by the results in Table 15 in this document.
    In Table 12 (in the Appendix) in the original paper, within “Basic UI” section, the data in the Majority Vote (MV) row was pasted incorrectly. The data presented in the original paper were mistakenly copied from the ZBS row under Advanced UI section in Table 11 (in the Appendix). This error did not affect our conclusion.
  3. (3)
    Additionally, the sentence on page 8 of the original paper,
    “We also noticed that in the simulations, accuracy improved for most aggregation methods as the number of MTurk workers increased, but MACE’s performance was notably inconsistent.”
    should be corrected as follows:
    “We also noticed that in the simulations, accuracy improved for most aggregation methods as the number of MTurk workers increased.”
    The last part in the original sentence referred to the unstable performance of the older tool used in early experiments, the Crowd-Kit implemented MACE method, as detailed in the sections 3.6.1 and 3.6.3 of the original paper. In the final version of the paper, we instead used Hovy’s MACE implementation, which we also described in the sections 3.6 of the original paper. Therefore, the last part of the sentence needs to be removed. This error did not impact the main conclusion or affect any other results presented in the original paper.
These corrections do not affect our overall conclusions. Our core findings remain valid.
Basic Interface
Method Acc P-Value 95% CI Acc P-Value 95% CI Acc P-Value 95% CI
MV .551 <.001 [.534, .569] .609 <.001 [.592, .626] .567 <.001 [.550, .584]
DawidSkene .675 <.001 [.659, .691] .737 <.001 [.722, .753] .712 <.001 [.696, .727]
OneCoin .797 .002 [.783, .811] .804 .002 [.790, .818] .800 .006 [.786, .814]
GLAD .734 <.001 [.719, .749] .756 <.001 [.741, .771] .731 <.001 [.715, .746]
M-MSR .531 <.001 [.513, .548] .630 <.001 [.614, .647] .590 <.001 [.572, .607]
MACE .811 .001 [.797, .824] .809 .001 [.795, .823] .809 <.001 [.796, .823]
Wawa .633 <.001 [.617, .650] .672 <.001 [.656, .689] .636 <.001 [.619, .653]
ZBS .677 <.001 [.661, .694] .712 <.001 [.696, .727] .675 <.001 [.659, .691]
Avg. Acc .676 - - .716 - - .690 - -
#workers 216 134 176
Table 14. Aggregation Accuracy Results of the Basic Interface integrated with GPT4 Group. Bold and underline highlight the highest score within the column and across the table, respectively. P-value is obtained by comparing with GPT-4 over the article-level accuracy. (**: p<0.01; ***: p<0.001. Paired t-test. N=200)
Eval Background Purpose Method Finding Other Acc Kappa
Label P R F1 P R F1 P R F1 P R F1 P R F1
Basic UI
MV .780 .387 .517 .194 .613 .294 .450 .710 .551 .854 .582 .692 1.000 .286 .444 .567 .402
Advances UI